# Enhanced entanglement scaling and area-law charge fluctuations in a non-Fermi liquid of composite fermions

### v1

@@ Abstract
In Fig. 1 we show the evaluated $S_{2}$ for bosonic and fermionic CFL states at fillings $\nu{=}1/m$, with $m{=}1,2,3,4$. Based on our free-fermion benchmarks, to minimize the finite-size effect, we harvest data points such that the subsystem area ranges from 1% to 30% of the total area, and use a fixed Fermi momentum of $k_{F}=\sqrt{2\nu}$. Our results for $\nu{=}1/2$ are consistent with Ref. [27] and scaling in Eq. (3), with the coefficient $\kappa{\approx}1.7\kappa_{\mathrm{w}}$ violating the Widom formula. By comparing the circular and square subregions, we find the violation to be independent of the subregion shape [29]. Furthermore, attaching a different number of vortices only has a weak effect on the slope $\kappa$. Moreover, LLL projection also has an almost negligible effect on $\kappa$, except for the bosonic $\nu{=}1$ CFL state. The stronger effect of LLL projection on $\nu{=}1$ is expected based on the single vortex attached [29]. Finally, while Fig. 1 is for the torus geometry, we find a similar enhancement of entanglement scaling for CFL states on the sphere [29].
@@
Charge fluctuations.—In systems with a U(1) symmetry, entanglement can be directly related to fluctuations of particle number, $\hat{N}_{A}$, in a subsystem:
@@
For example, in non-interacting 2D Fermi liquids, $S_{2}$ and $\Delta^{2}N_{A}$ differ only by a multiplicative constant [44, 45]. Similarly, in a NFL, charge fluctuations are expected to closely track entanglement entropy [46]. Recently, the scaling of charge cumulants or “disorder operators” was indeed shown to exhibit a logarithmic violation of the area law in NFLs at a quantum critical point [47]. Since $\Delta^{2}N_{A}$ can be computed with a similar Monte Carlo method and system sizes as $S_{2}$, we can leverage the better analytical understanding of charge fluctuations to support our previous entanglement results.
Remarkably, Fig. 2 shows that $\Delta^{2}N_{A}$ in CFL states obeys an area law, irrespective of LLL projection and in stark contrast with ordinary Fermi liquids. The electrons in a CFL have been argued to exhibit an effective finite correlation length, leading to an exponentially suppressed tunneling density of states [48, 49, 50, 51] and a Coulomb gap observed in experiment [52, 53, 54, 55, 56]. This implies that charge fluctuations of the physical electrons forming the CFL are gapped and should follow an area law [31, 57]. This striking property of the CFL state, to the best of our knowledge, has not previously been demonstrated for the CFL wave function.
In spatially isotropic systems, the behavior of $\Delta^{2}N_{A}$ is governed by the static structure factor [33]. For CFL states, the LLL-projected structure factor takes the form $s_{q}{\sim}q^{3}\ln(1/q)$ in the long-wavelength limit [2, 11], with inter-Landau level gapped modes contributing an additional term ${\sim}q^{2}$. The former enforces an area law scaling, while the latter adds a “super-universal” constant term in the presence of sharp corners, e.g., for the square subsystem in Fig. 1(a). The general form is
@@
where $a$ is a non-universal coefficient, and $b(\theta)$ is a constant contribution due to a bipartition containing a single corner with an opening angle $\theta$ [31]. From the intercept of $\Delta^{2}N_{A}$ vs. $\lambda$, we determined $b(\pi/2)$ in Table 1, which is in good agreement with Eq. (5). Note that while projection places the wave functions of Eq. (1) completely inside the LLL, we do not project the density operator, and therefore the associated gapped modes are not removed, hence both projected and unprojected wave functions exhibit approximately the same corner contribution.
@@
Figure 2: Charge fluctuations across the bipartition for CFL states at different filling factors. In contrast to free fermions (solid black line), the charge fluctuations in CFL states are strongly suppressed and obey area-law scaling (plots on a linear $\lambda$ scale are shown in SM [29] and the colored dashed lines represent fits to the area law). LLL projection alters the area law coefficient but preserves the universal corner contribution. Table 1 shows the extracted corner contributions, which are proportional to the filling.
@@
Modulus and sign structure.—While Fig. 1 is strongly suggestive of $\kappa$ exceeding the Widom value, it is important to understand where this enhancement is structurally encoded in a CFL wave function and whether or not it should be interpreted as a finite-size effect [28]. An idea first put forward in Ref. [58] in the context of gapped spin liquids, is to decompose the total entropy in a way inspired by SWAP:
$S_{2}(\Psi)=S_{2}(|\Psi|)+S_{2,\mathrm{sign}},$ (6)
@@
where $S_{2}(|\Psi|)$ is the entropy of the absolute value of the wave function, while $S_{2,\mathrm{sign}}$ stems from an interplay between the absolute value and the sign structure of $\Psi$ [29].
@@
The hope behind the decomposition of Eq. (6) is that the sign structure of the wave function carries the leading contribution to the entanglement as $N{\to}\infty$. While there are examples of sign wave functions with provable large entanglement [59, 60], the scaling of different terms with $N$ in Eq. (6) is not rigorously understood in general. Indeed, even for free fermions, the scaling of $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ are not known analytically. Our Monte Carlo analysis for free fermions in SM [29] is consistent with $S_{2,\mathrm{sign}}$ carrying the dominant $\lambda\ln\lambda$ dependence, however both $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ suffer from much more pronounced finite-size effects compared to the total entropy $S_{2}$. Consequently, the decomposition given in Eq. (6) obscures the interpretation of the data as it replaces a tractable free-fermion problem with an effectively “many-body” object that carries non-trivial finite-size dependence due to the mixing of subleading [$\mathcal{O}(\lambda)$ and $\mathcal{O}(\ln\lambda)$] terms.
@@
Figure 3: Sign entropy for projected (a) and unprojected (b) CFL states for $N=37$ particles on a torus. For comparison, we also show the result for free fermions, with the same data plotted in (a) and (b). Unprojected CFL states appear to behave similarly to free fermions, while projected wave functions show an enhanced slope. This is revealed by plotting the difference of $S_{2,\mathrm{sign}}$ for the $\nu{=}1/2$ CFL relative to free fermions in panels (c)-(d): the difference grows in the projected case (c), while it decays towards zero in the unprojected case (d). Insets in (c)-(d) show the ratio of CFL sign entropy to that of free fermions, which shows qualitatively similar behavior.
@@
Similar issues are observed in the sign entropy of projected and unprojected CFL states, shown in Fig. 3(a)-(b). Unlike the full entropy, which is largely insensitive to LLL projection, the sign entropies are significantly altered by it, although they still appear to capture the logarithmic violation of the area law. In particular, $S_{2,\mathrm{sign}}$ of the unprojected CFL states tracks the sign entropy of free fermions, which is revealed more clearly by plotting their difference in Fig. 3(d), while the ratio is shown in the inset. By contrast, within the same range of system sizes, the difference $S_{2,\mathrm{sign}}{-}S_{2,\mathrm{sign}}^{\mathrm{free}}$ increases with $\lambda$ for the projected CFL states in Fig. 3(c), consistent with an enhancement.
@@
As the physical content and scaling properties of the sign entropy are unclear, one should not view the results in Fig. 3 as more compelling than the full entropy in Fig. 1. Nevertheless, if one assumes that $S_{2,\mathrm{sign}}$ may carry the leading $\lambda\ln\lambda$ dependence, it is interesting to ponder two scenarios that would make Figs. 1 and 3 consistent. The first possibility is that there may be an intrinsic difference between projected and unprojected CFL states, such that only the former exhibits an enhanced entanglement growth. While this is counterintuitive, LLL projection could alter the coefficient of entanglement scaling, unlike other universal topological properties. The second option, proposed by Ref. [28], is to conclude there is no enhancement and place trust in Figs. 3(b)-(d) above all others. This would require the lines in Figs. 3(a)-(c) to ultimately undergo a downturn, with a corresponding reduction of slope in Fig. 1. Based on the existing data, we estimate that observing this would require system sizes $N{\gtrsim}300$, which are significantly beyond current computational capability.
The similarity of our unprojected results in Fig. 3(b) with Ref. [28] is attributed to the lack of an explicit band projection in the lattice realization of the CFL state. By contrast, our LLL-projected results in Fig. 1 show a visible departure from Ref. [28], e.g., the entropy of the bosonic $\nu{=}1$ CFL is similar to that of the fermionic $\nu{=}1/2$ CFL state after LLL projection, unlike their lattice analogs in Ref. [28]. These differences point to a subtle role of LLL projection, precluding direct comparisons between continuum and lattice versions of the CFL state, and potentially implying that entanglement scaling may depend on the microscopic details of flux attachment. A related question is whether the Chern band projection has a similar effect on the entanglement of CFL in zero magnetic field [6, 7], or whether the latter also depends on band geometry and Chern number.

### v2

@@ II.1 Torus
Our results for $\nu=1/2$ in Fig. 1(a) are consistent with Ref. [27] and scaling in Eq. (4), with the coefficient $\kappa\approx 1.7\kappa_{\mathrm{w}}$ significantly violating the Widom formula. By comparing the circular and square subregions, we find the violation to be independent of the subregion shape, see Appendix D. Another thing to note is that attaching a different number of vortices only has a weak effect on the slope $\kappa$, as seen from the parallel lines in Fig. 1(a). Moreover, LLL projection also has an almost negligible effect on $\kappa$, except for the bosonic $\nu=1$ CFL state, as illustrated in Fig. 1(c). The stronger effect of LLL projection on $\nu=1$ is expected based on the single vortex attached [53]. Next, we show that similar results are obtained for CFL states on the sphere.
II.2 Sphere
The unprojected CFL wave function on the sphere is of the form in Eq. (1), where the single-particle orbitals are the standard spherical harmonics, $\chi_{n}(\mathbf{r}_{j})=Y_{L_{n}M_{n}}(\theta_{j},\varphi_{j})$, where $L$, $M$ stand for the angular momentum and its $z$-component, respectively, while $\theta_{j}$ and $\phi_{j}$ are the polar and azimuthal angles describing the position of $j$th particle in terms of spherical angle $\mathbf{\Omega}_{j}$. The spherical Laughlin-Jastrow wave function has a compact expression
$\psi^{\mathrm{L}}_{m}=\prod_{i<j}(u_{i}v_{j}-v_{i}u_{j})^{m},$ (8)
in spherical spinor coordinates [54],
$u_{j}=\sin(\theta_{j}/2)e^{i\phi_{j}/2},\;\;\;v_{j}=\cos(\theta_{j}/2)e^{-i\phi_{j}/2}.$ (9)
Note that the Widom formula for free fermions can also be appropriately modified for the sphere geometry, as shown in Appendix E.
Before presenting the results, we briefly explain how to perform the LLL projection on the sphere. Similar to the torus, while Eq. (8) describes a proper LLL state, the spherical harmonics have components in arbitrary LLs. Let us first treat the state at $\nu=1/2$, with two vortices attached i.e., $\Psi^{\rm CFL}_{1/2}{=}\mathcal{P}_{\rm LLL}\prod_{j{<}k}(u_{j}v_{k}{-}u_{k}v_{j})^{2}\Phi_{\rm FS}$, where $\Phi_{\rm FS}$ is the filled-shell wave function on the sphere at zero field that is analogous to that of a Fermi sea [26]. The projected state can be evaluated as a Slater determinant of CF wave functions $Y^{CF}_{0LM}$ which, similarly to the torus geometry, depend on the coordinates of all particles through the Jastrow factors $J_{j}{=}\prod_{k{\neq}j}(u_{j}v_{k}{-}u_{k}v_{j})$. We quote the final expression here:
$Y^{CF}_{0LM}(\mathbf{\Omega}_{j})$ $=$ $N_{0,L,M}(-1)^{L-M}\frac{(2Q+1)!}{(2Q+L+1)!}u_{j}^{M}v_{j}^{-M}$ (10)
$\sum_{s=0}^{L}(-1)^{s}\begin{pmatrix}L\\ s\end{pmatrix}\begin{pmatrix}L\\ L-M-s\end{pmatrix}u_{j}^{s}v_{j}^{L-s}$
$\bigg{[}\bigg{(}\frac{\partial}{\partial u_{j}}\bigg{)}^{s}\bigg{(}\frac{\partial}{\partial v_{j}}\bigg{)}^{L-s}J_{j}\bigg{]}\,,$
where $N_{0,L,M}$ is a normalization factor, and the flux for the physical electrons is $2Q=2(N{-}1)$. Given that we only have one power of Jastrow, the derivatives can be evaluated as follows [9]:
$\bigg{(}\frac{\partial}{\partial u_{j}}\bigg{)}^{s}\bigg{(}\frac{\partial}{\partial v_{j}}\bigg{)}^{L-s}J_{j}$ (11)
$=$ $J_{j}\sum\prod_{i=0}^{s}\frac{v_{k_{i}}}{u_{j}v_{k_{i}}-u_{k_{i}}v_{j}}\prod_{i=0}^{L-s}\frac{-u_{l_{i}}}{u_{j}v_{l_{i}}-u_{l_{i}}v_{j}}\,,$
where the sum runs over all possibilities of choosing $s$ particles $\{k_{i}\}$ and $L{-}s$ particles $\{l_{i}\}$, all distinct. The Jastrow factors $J_{j}$ can be factored out of the determinant post-projection and their product is the square of the complete Jastrow factor, i.e., $\prod_{j}J_{j}=\prod_{i<k}(u_{i}v_{k}-u_{k}v_{i})^{2}\equiv\psi^{\rm L}_{2}$.
The above construction produces CFL states at general fillings. For the fermionic state at $\nu=1/4$, we choose to only project with two Jastrow factors, leaving the other two outside – this is similar to the $\bm{\alpha}=(4,0)$ state on the torus. For the bosonic states at $\nu=1$ and $\nu=1/3$, we need to divide or multiply by a single Jastrow factor respectively [55, 30], again mirroring the construction on the torus. It is known that the microscopic CFL wave function is not very sensitive to the number of Jastrows placed inside the LLL-projection operator [56, 57]. Note that $\nu=1$ is a special case as as the hardcore constraint is not enforced properly, which can result in instability after projection. To fix this, we implicitly enforce a non-zero hardcore radius in our Monte Carlo simulation.
Figure 2: $S_{2}$ Rényi entropy of projected and unprojected CFL states on the sphere at fillings $\nu=1/m$, with $m=1,2,3,4$. The projected data is for the system size $N=49$, while the unprojected data is for $N=64$. The continuous black line is free-fermion data obtained by exact diagonalization. All filling factors show an enhanced slope compared to the Widom value, and $\nu=1$ shows an increase after projection, similar to the torus geometry.
Fig. 2 demonstrates the similarity of CFL results on the sphere with those presented previously for the torus geometry. This consistency highlights the robustness of our results. Furthermore, for the sequence of system sizes $N=n^{2}$ we consider on the sphere, the CFL has a uniform angular momentum $L=0$ state, which removes some of the ambiguity associated with the definition of the Fermi momentum on the torus.
III Charge fluctuations
@@
In systems with a U(1) symmetry, entanglement can be directly related to the fluctuations of particle number, $\hat{N}_{A}$, measured on a subsystem:
@@
For example, in non-interacting 2D Fermi liquids, $S_{2}$ and $\Delta^{2}N_{A}$ differ only by a multiplicative constant [58, 59]. Similarly, in a NFL, charge fluctuations are expected to closely track entanglement entropy [60]. Recently, the scaling of charge cumulants or “disorder operators” was indeed shown to exhibit a logarithmic violation of the area law in NFLs at a quantum critical point [61]. Since $\Delta^{2}N_{A}$ can be computed with a similar Monte Carlo method and the same system sizes as $S_{2}$, we can leverage the better analytical understanding of charge fluctuations to support our previous entanglement results.
Remarkably, Fig. 3 shows that $\Delta^{2}N_{A}$ in CFL states, with and without LLL projection, obeys an area law – in stark contrast with ordinary Fermi liquids. This can be understood from the fact that the electrons in the CFL exhibit an effectively finite correlation length, leading to an exponentially suppressed tunneling density of states [62, 63, 64, 65] and a Coulomb gap observed in experiment [66, 67, 68, 69, 70]. This implies that charge fluctuations of the physical electrons forming the CFL are gapped and follow an area law [32, 71]. This striking property of the microscopic CFL wave function, to the best of our knowledge, has been demonstrated for the first time in this work.
In spatially isotropic systems, the static structure factor $s_{q}$ governs the behavior of $\Delta^{2}N_{A}$ [34]. For CFL states, the LLL-projected structure factor takes the form $\bar{s}_{q}\sim q^{3}\ln(1/q)$ in the long-wavelength limit [2, 11], with inter-Landau level gapped modes contributing an additional term $\sim q^{2}$ to $s_{q}$. The former enforces an area law scaling, while the latter adds a “super-universal” constant term in the presence of sharp corners. The general form is
@@
where $a$ is a non-universal coefficient, and $b(\theta)$ is a constant contribution due to a bipartition containing a single corner with an opening angle $\theta$ [32]. In Fig. 3, for simplicity, we consider a square subsystem.
From the intercept of $\Delta^{2}N_{A}$ vs. $\lambda$, we determined $b(\pi/2)$ in Table 1, which is in good agreement with Eq. (13). Note that while projection places the wave functions of Eq. (1) completely inside the LLL, we do not project the density operator, and therefore the associated gapped modes are not removed, hence both projected and unprojected wave functions exhibit approximately the same corner contribution.
@@
Figure 3: Charge fluctuations across the square-shaped bipartition for CFL states at different filling factors. In contrast to free fermions (solid black line), the charge fluctuations in CFL states are strongly suppressed and obey area-law scaling. The top row shows the fluctuations divided by $\lambda$ and plotted as a function of $\ln\lambda$, while the bottom row shows the raw value plotted on a linear $\lambda$ scale. Insets to the bottom plots show fits to the area law (colored dashed lines). LLL projection alters the area law coefficient but preserves the universal corner contribution. Table 1 shows the extracted corner contributions, which are proportional to the filling factor.
Figure 4: Comparison of the full $S_{2}$ Rényi entropy and its SWAP decomposition, Eq. (14), for free fermions on the sphere. (a) Total $S_{2}$ entropy, computed via Monte Carlo (markers), while exact diagonalization result (solid line) is for a large system size $N=625$. All system sizes steadily approach the Widom slope $\kappa_{\mathrm{w}}{=}0.25$ (dashed line). (b): $S_{2,\mathrm{p}}$ and $S_{2,\mathrm{mod}}$ that sum up to the absolute value entropy $S_{2}(|\Psi|)$. $S_{2,\mathrm{p}}$ appears to scale logarithmically with the subsystem size, while $S_{2,\mathrm{mod}}$ follows an area-law. Black lines are $\lambda\to\infty$ extrapolations. (c) The absolute value entropy $S_{2}(|\Psi|)=S_{2,p}+S_{2,\mathrm{mod}}$. The black line is the sum of the extrapolations in (b), which eventually settles to an area law for very large systems. However, unlike the individual decompositions in (b), their sum here appears far from the asymptotic scaling regime. (d) The sign entropy $S_{2,\mathrm{sign}}$. The dashed line is the Widom slope, while the continuous black line is the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$. This highlights the difficulty in extracting the Widom slope solely from the sign term: the continuous and dashed lines only become parallel for $\ln\lambda\gtrsim 4$, effectively requiring much larger system sizes to infer the correct scaling compared to the full entropy in (a).
IV Modulus and sign structure
@@
While Fig. 1 strongly suggests that $\kappa$ exceeds the Widom value, it is important to understand where this enhancement is structurally encoded in a CFL wave function and whether or not it should be interpreted as a finite-size effect [28]. An idea first put forward in Refs. [72, 73] is to decompose the total entropy in a way inspired by SWAP:
$S_{2}(\Psi)=S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}++S_{2,\mathrm{sign}}.$ (14)
@@
The first term represents the probability that the two copies are “swappable”, while the last two terms are the “mod” and “sign” contributions – see Appendix A. Adding the first two contributions together yields $S_{2}(|\Psi|)$ – the entropy of the absolute value of the wave function, while $S_{2,\mathrm{sign}}$ stems from an interplay between the absolute value and sign structure of $\Psi$.
@@
The hope behind the decomposition of Eq. (14) is that the sign structure of the wave function carries the leading contribution to the entanglement as $N{\to}\infty$. While there are examples of sign wave functions with provable large entanglement [74, 75], there are also instances where the mod entropy is not dominated by that of the sign [76]. Thus, the scaling of different terms with $N$ in Eq. (14) is not rigorously understood in general. In fact, even for free fermions, the scalings of $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ are not known analytically. Our Monte Carlo analysis for free fermions presented below is consistent with $S_{2,\mathrm{sign}}$ carrying the dominant $\lambda\ln\lambda$ dependence, however both $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ suffer from much more pronounced finite-size effects compared to the total entropy $S_{2}$. Consequently, we will argue that the decomposition in Eq. (14) obscures the interpretation of the data as it replaces a tractable free-fermion problem with an effectively “many-body” object that carries non-trivial finite-size dependence due to the mixing of subleading [$\mathcal{O}(\lambda)$ and $\mathcal{O}(\ln\lambda)$] terms. Below we discuss in detail the entropy decompositions for free fermions and then apply the same considerations to the CFL case.
IV.1 Mod and sign entropy for free fermions
The scaling of different contributions to the entropy in Eq. (14) is not rigorously known, even for free fermions where the full entropy has an established analytic expression. Fig. 4(a) contrasts the full $S_{2}$ entropy for free fermions against its decompositions in Eq. (14) obtained via Monte Carlo, Fig. 4(b)-(d). While the full entropy can be computed for large free-fermion systems containing thousands of particles, the evaluation of mod-sign entropy contributions, to the best of our knowledge, can only be done with Monte Carlo. This limits the accessible system sizes to $N\lesssim 100$, which is similar to the range for the CFL states.
Within the range $N\lesssim 100$, we observe the full entropy $S_{2}$ steadily converging to the Widom slope in Fig. 4(a). Figs. 4(b)-(c) illustrate the importance of considering the $\mathrm{p}$ and $\mathrm{mod}$ contributions to the entropy separately. They each appear to converge nicely to the respective scalings, $S_{2,p}\sim\mathcal{O}(\ln\lambda)$ and $S_{2,\mathrm{mod}}\sim\mathcal{O}(\lambda)$. However, if we add them together to make up $S_{2}(|\Psi|)$, Fig. 4(c), the latter converges non-monotonically when plotted as $S_{2}(|\Psi|)/\lambda$ vs $\ln\lambda$. If we estimate the slope of $S_{2}(|\Psi|)$ by adding the fits, we obtain the solid line in Fig. 4(c), which suggests that our system sizes are not large enough to see the expected area-law scaling $S_{2}(|\Psi|)\sim\lambda$ in the thermodynamic limit. This has an adverse impact on the scaling of the sign entropy in Fig. 4(d), which overshoots the Widom slope to compensate for the downturn in $S_{2}(|\Psi|)$. By plotting the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$ using a solid line in Fig. 4(d), we see that it might only approach the Widom scaling in much larger system sizes ($\ln\lambda\gtrsim 4$). At the same time, the full entropy in Fig. 4(a) tracks the Widom scaling much more closely even in smaller system sizes. As the range of system sizes in this example is comparable to the CFL case, the usefulness of mod and sign decompositions over the full entropy is questionable.
IV.2 Mod and sign entropy for CFL
@@
Figure 5: Sign entropy for projected (a) and unprojected (b) CFL states for $N=37$ particles on a torus. For comparison, we also show the result for free fermions, with the same data plotted in (a) and (b). Unprojected CFL states behave similarly to free fermions, while projected wave functions show an enhanced slope. This is revealed by plotting the difference of $S_{2,\mathrm{sign}}$ for the $\nu=1/2$ CFL relative to free fermions in panels (c)-(d): the difference grows in the projected case (c), while it decays towards zero in the unprojected case (d). Insets in (c)-(d) show the ratio of CFL sign entropy to that of free fermions, which shows qualitatively similar behavior.
@@
Similar issues are observed in the sign entropy of projected and unprojected CFL states, Fig. 5(a)-(b). Unlike the full entropy, which is largely insensitive to the LLL projection, the sign entropies are significantly altered by it, although they still appear to capture the logarithmic violation of the area law. In particular, $S_{2,\mathrm{sign}}$ of the unprojected CFL states tracks the sign entropy of free fermions, which is revealed more clearly by plotting their difference in Fig. 5(d), while the ratio is shown in the inset. By contrast, within the same range of system sizes, the difference $S_{2,\mathrm{sign}}{-}S_{2,\mathrm{sign}}^{\mathrm{free}}$ increases with $\lambda$ for the projected CFL states in Fig. 5(c), consistent with an enhancement.
@@
As the physical content and scaling properties of the sign entropy are unclear, one should not view the results in Fig. 5 as more compelling than the full entropy in Fig. 1. Nevertheless, if one assumes that $S_{2,\mathrm{sign}}$ may carry the full $\lambda\ln\lambda$ dependence, it is interesting to ponder two scenarios that would make Figs. 1 and 5 consistent. The first possibility is that there may be an intrinsic difference between projected and unprojected CFL states, such that only the former exhibits an enhanced entanglement growth. While this is counterintuitive, LLL projection could alter the coefficient of entanglement scaling, unlike other universal topological properties. The second option, proposed by Ref. [28], is to conclude there is no enhancement and place trust in Figs. 5(b)-(d) above all others. This would require the lines in Figs. 5(a)-(c) to ultimately undergo a downturn, with a corresponding reduction of slope in Fig. 1. Based on the existing data, we estimate that observing this would require system sizes $N{\gtrsim}300$, which are significantly beyond current computational capability.
V Conclusions and discussion